OpenStudy (anonymous):

find the probability the first three cards chosen are a queen and 2 diamonds.

OpenStudy (anonymous):

I need help with stats

OpenStudy (anonymous):

so there's 52 cards in a deck, 4 queens in all(1 in each club) and i think 13 diamonds (1-10, the king and queen, and the jack), so take the combined probability of the queens and diamonds 17/52 and multiply it by 3/52

OpenStudy (anonymous):

So do I multiply since its "and"

OpenStudy (anonymous):

multiply the combined probability of the total queens and diamonds in a deck by the probability of the cards at hand which is 3 out of 52.

OpenStudy (anonymous):

Oh alright, thank you very much

OpenStudy (anonymous):

and with that answer too(i forgot 2/3 were diamonds) multiply the ratio of two stacks of diamonds which is 26/52=1/2 by that previous result(17/52*3/52) and do the same for 1 queen 1/52

OpenStudy (anonymous):

oh that's what was missing

OpenStudy (anonymous):

yeah i had to clean that up ;)

OpenStudy (anonymous):

thx buddy

OpenStudy (kropot72):

If we assume that the chosen queen can not be the queen of diamonds, the queen of diamonds needs to excluded from the choice of queen (giving 3 queens to choose from) and also excluded from the choice of the 2 diamonds (avoiding getting 2 queens). The required probability is therefore given by: \[\large P(1\ queen,\ 2\ diamonds)=\frac{3C1\times 12C2}{52C3}=\frac{3\times12\times11\times3\times2}{2\times52\times51\times50}=you\ can\ calculate\]

OpenStudy (anonymous):

hmmm, didn't know i had to use the combination rule

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome. The problem is sampling without replacement of the multivariate type. "Find the probability the first three drinks chosen are diet sprite and 2 waters (in that order)" Please post as a new question. Note:The populations of the types of drink needs to be included to make a solution possible.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry for that